OpenStudy (anonymous):

Solve the variable 6x - 6 = 7x - 5

OpenStudy (praxer):

6x-6=7x-5 =>6x-7x=-5+6 =>-x=1 =>x=-1

OpenStudy (praxer):

did you understand this....? ?......:)

OpenStudy (anonymous):

Yes, thank you. I had the same answer but was afraid I was wrong lol Thanks

OpenStudy (praxer):

$$\huge \frak Welcome \ddot\smile $$